OpenStudy (anonymous):

Find the derivative... (t)/((t^2+t^9)(t^5+1))

OpenStudy (aum):

u = t u' = 1 v = (t^2+t^9) * (t^5+1) v' = ? d/dt (u/v) = (v * u' - u * v') / v^2